# Testing Monotonicity in a Finite Population

### v1

@@ 2. Case 2
Now, let $\delta^{*}(Y)=1\{Y\in S^{Y}_{\theta_{1}}\}$. Since $S^{Y}_{\theta_{1}}\not\subseteq S^{Y}_{\theta_{0}}$ for all $\theta_{0}\in\Theta_{0}$, it follows that $E_{\theta_{0}}[\delta^{*}(Y)]<1$ for all $\theta_{0}\in\Theta_{0}$. Since $\Theta_{0}$ is finite, we thus obtain that $\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta^{*}(Y)]<1$. However, by construction $E_{\theta_{1}}[\delta^{*}(Y)]=1$. Hence, $\delta^{*}(Y)$ controls size at level $\alpha^{*}=\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta^{*}(Y)]<1$ and has power of 1 against the alternative $\theta_{1}$. If $\alpha^{*}<\alpha$, then the proof is complete by setting $\delta(Y)=\delta^{*}(Y)$ (and thus we have a non-randomized test). If $\alpha^{*}>\alpha$, set $\delta(Y)=\frac{\alpha}{\alpha^{*}}\delta^{*}(Y)$. Then by construction $\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta(Y)]=\alpha$ and $E_{\theta_{1}}[\delta(Y)]=\alpha^{*}>\alpha$. ∎
@@
Let $\pi(t;\theta)$ be the PMF for $\mathrm{Multinomial}(n,\theta/n)$. Then
$w_{n}(t;\theta)=\frac{\pi(t;\theta)}{\sum_{t\in\Theta_{1}}\pi(t;\theta)}\mathbbm{1}(t\in\Theta_{1}).$
Define $g_{\theta}(y)=\sum_{t\in\Theta_{0}\cup\Theta_{1}}f_{t}(y)\pi(t;\theta)$. By Proposition˜3.1, there exists some $p\in\mathcal{P}_{0}$ such that $g_{\theta}(y)=g_{p}(y)$. We thus have the following expansion of WAP,
$\mathrm{WAP}(\delta;\theta)$ $=\sum_{t\in\Theta_{0}\cup\Theta_{1}}\sum_{y}\delta(y)f_{t}(y)\frac{\pi(t;\theta)}{\sum_{t\in\Theta_{1}}\pi(t;\theta)}\mathbbm{1}(t\in\Theta_{1})$
$=\underbrace{\left({\sum_{t\in\Theta_{1}}\pi(t;\theta)}\right)^{-1}}_{\equiv Z_{1}^{-1}}\sum_{y}\delta(y)\left[{\sum_{t\in\Theta_{0}\cup\Theta_{1}}f_{t}(y)\pi(t;\theta)-\sum_{t\in\Theta_{0}}f_{t}(y)\pi(t;\theta)}\right]$
$=\frac{1}{Z_{1}}\left[{E_{g_{\theta}}[\delta(Y)]-\sum_{t\in\Theta_{0}}E_{t}[\delta(Y)]\pi(t;\theta)}\right]$
@@
where recall $v=\frac{1}{n}\min\{\theta_{c},\theta_{d}\}=\min\{q_{c},q_{d}\}$. This proves the first inequality. The second inequality follows from examining that $v\geq 1/n$ and that
@@
Let $\delta_{0}(Y)=\alpha\cdot 1\{(Y_{T},Y_{U})\not\in\{(n_{1},0),(0,n_{0})\}\}$ be the test that rejects with probability 0 if $Y$ is either $(n_{1},0)$ or $(0,n_{0})$, and rejects with probability $\alpha$ otherwise. Let $S^{Y}_{\theta}$ be the support of $Y$ under parameter $\theta$. Note that by construction, for any $\theta_{0}\in\Theta_{0}$, we have that
@@
$E_{\theta_{0}}[\delta_{0}(Y)]=\alpha\text{ if }E_{\theta_{0}}[\delta_{1}(Y)]=0.$
@@
it suffices to show that
@@
Finally, we show that there exists $\theta_{1}$ such that $E_{\theta_{1}}[\delta(Y)]>\alpha$. Since we showed above that $E_{\theta_{1}}[\delta_{0}(Y)]=\alpha$ for all $\theta_{1}\in\Theta_{1}$, it suffices to show that there exists some $\theta_{1}\in\Theta_{1}$ such that $E_{\theta_{1}}[\delta_{1}(Y)]>0$, or equivalently $P_{\theta_{1}}(Y=(n_{1}-1,1))>0$. However, if $\theta_{1}$ corresponds to $(n_{at},n_{nt},n_{d},n_{c})=(0,0,1,n-1)$, then $Y=(n_{1}-1,1)$ obtains when the defier is assigned to control. ∎
@@
$E_{\theta}[\delta(Y)]\leq\alpha(1+3.125n^{1.5}(1-\epsilon)^{n})\to\alpha.$
@@
$\pi(\theta_{0}\mid Y=y)=1-\pi(\theta_{1}\mid Y=y)=\frac{f_{\theta_{0}}(y)}{f_{\theta_{1}}(y)+f_{\theta_{0}}(y)}.$
Since $f_{\theta_{0}}\neq f_{\theta_{1}}$ for some $y$ in the support of one of $f_{\theta_{0}}(\cdot)$ and $f_{\theta_{1}}(\cdot)$, $\pi(\theta_{0}\mid Y=y)\neq 0.5=\pi(\theta_{0})$ with positive probability. Since $\theta_{0}\in\Theta_{0}$ is the only point of support under $\pi$, which concludes the proof of the first part.

### v2

@@ 2. Case 2
Now, let $\delta^{*}(Y)=1\{Y\in S^{Y}_{\theta_{1}}\}$. Since $S^{Y}_{\theta_{0}}\not\subseteq S^{Y}_{\theta_{1}}$ for all $\theta_{0}\in\Theta_{0}$, it follows that $E_{\theta_{0}}[\delta^{*}(Y)]<1$ for all $\theta_{0}\in\Theta_{0}$. Since $\Theta_{0}$ is finite, we thus obtain that $\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta^{*}(Y)]<1$. However, by construction $E_{\theta_{1}}[\delta^{*}(Y)]=1$. Hence, $\delta^{*}(Y)$ controls size at level $\alpha^{*}=\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta^{*}(Y)]<1$ and has power of 1 against the alternative $\theta_{1}$. If $\alpha^{*}\leq\alpha$, then the proof is complete by setting $\delta(Y)=\delta^{*}(Y)$ (and thus we have a non-randomized test). If $\alpha^{*}>\alpha$, set $\delta(Y)=\frac{\alpha}{\alpha^{*}}\delta^{*}(Y)$. Then by construction $\sup_{\theta_{0}\in\Theta_{0}}E_{\theta_{0}}[\delta(Y)]=\alpha$ and $E_{\theta_{1}}[\delta(Y)]=\frac{\alpha}{\alpha^{*}}>\alpha$. ∎
Lemma A.1.
Suppose $n$ is even. Then there exist priors $\pi_{A}$ and $\pi_{B}$ on $\theta$ such that $\pi_{A}(\theta\in\Theta_{0})=1$, $\pi_{B}(\theta\in\Theta_{0})=0$, and $\pi_{A}(Y=y)=\pi_{B}(Y=y)$ for all $y\in\{0,\ldots,n_{1}\}\times\{0,\ldots,n_{0}\}$.
Proof.
Define $\tilde{\pi}_{A}=\tilde{\pi}_{(0.5,0.5,0,0)}$ and $\tilde{\pi}_{B}=\tilde{\pi}_{(0,0,0.5,0.5)}$. The prior $\tilde{\pi}_{A}$ arises from assuming that the units in the finite population were sampled i.i.d. from a superpopulation in which half of individuals are always-takers and the other half are never-takers. The prior $\tilde{\pi}_{B}$ arises analogously if the superpopulation is half compliers and half defiers. As argued in the proof to Proposition˜3.1, the marginal distribution of $Y$ after (i) sampling $\theta\sim\mathrm{Multinomial}(1,p)$ and (ii) generating $Y\sim f_{\theta}$ depends on $p$ only through $p_{at}+p_{c}$ and $p_{at}+p_{d}$. It follows that the two priors imply the same unconditional distribution for $Y$, $\tilde{\pi}_{A}(Y=y)=\tilde{\pi}_{B}(Y=y)$ for all $y$.
Now, consider type counts $\theta_{0}(n_{at})=(n_{at},n-n_{at},0,0)$, which have positive numbers of only always-takers and never-takers. Observe that the support of $Y$ under $\theta_{0}(n_{at})$ is
$S^{Y}_{\theta_{0}(n_{at})}=\{(m,n_{at}-m)\,:\,m=\max\{0,n_{at}-n_{0}\},\ldots,\min\{n_{1},n_{at}\}\},$ (2)
where $(m,n_{at}-m)$ is the realization of $Y$ if $m$ of the always-takers are selected for treatment. Hence the support points for $Y$ under $\theta_{0}$ lie on a line with slope of 1-1 and intercept $n_{at}$. Letting $S^{Y}_{\theta_{0}}$ denote the support of $Y$ under $\theta_{0}$, we see that the $S^{Y}_{\theta_{0}}$ are disjoint for all $\theta_{0}$ to which $\tilde{\pi}_{A}$ assigns positive support:
$S^{Y}_{\theta_{0}(n_{at})}\cap S^{Y}_{\theta_{0}(n_{at}^{\prime})}=\varnothing\text{ for $n_{at}\neq n_{at}^{\prime}$}.$
$\tilde{\pi}_{A}(Y=y\mid\theta\in\tilde{\Theta}_{0})=\tilde{\pi}_{A}(Y=y\mid Y\in S_{\tilde{\Theta}_{0}}),$
Similarly, let $\theta_{1}(n_{c})=(0,0,n-n_{c},n_{c})$, which has only compliers and defiers. The support points of $Y$ under $\theta_{1}$ take the form
$S^{Y}_{\theta_{1}(n_{c})}=\{(m,m+n_{0}-n_{c})\,:\,m=\max\{0,n_{c}-n_{0}\},\ldots,\min\{n_{1},n_{c}\}\},$ (3)
where $(m,n_{0}-(n_{c}-m))=(m,m+n_{0}-n_{c})$ corresponds to the realized outcomes when $m$ compliers are selected for treatment. These points lie on an upward sloping line with slope 1 and intercept $n_{0}-n_{c}$. Letting $S^{Y}_{\theta_{1}}$ denote the support of $Y$ under $\theta_{1}$, we again see that the $S^{Y}_{\theta_{1}}$ are disjoint. Hence, for any set $\tilde{\Theta}_{1}$ of the form $\{\theta_{1}(n_{c,1}),\ldots,\theta_{1}(n_{c,K})\}$, we have that
$\tilde{\pi}_{B}(Y=y\mid\theta\in\tilde{\Theta}_{1})=\tilde{\pi}_{B}(Y=y\mid Y\in S_{\tilde{\Theta}_{1}}),$
Now, let
$\tilde{\Theta}_{0}\coloneqq\bigcup_{\begin{subarray}{c}0\leq n_{at}\leq n\\ n_{at}\equiv n_{1}+1\,(\mathrm{mod}\ 2)\end{subarray}}\{\theta_{0}(n_{at})\}\subset\Theta_{0}$ (4)
and
$\tilde{\Theta}_{1}\coloneqq\bigcup_{\begin{subarray}{c}1\leq n_{c}\leq n-1\\ n_{c}\ \mathrm{odd}\end{subarray}}\{\theta_{1}(n_{c})\}\subset\Theta_{1}.$ (5)
Lemma˜A.2 shows formally that
$S_{\tilde{\Theta}_{1}}=S_{\tilde{\Theta}_{0}}.$
An intuitive illustration of the proof is shown in Figure˜2 for the setting where $n_{1}=n_{2}=2$, in which case $S_{\tilde{\Theta}_{1}}$ corresponds to the union of the points on the two upward-sloping lines in orange, and $S_{\tilde{\Theta}_{0}}$ corresponds the unions of the points on the two downward-sloping lines in blue.
[…]
Figure 2. Illustration of the sets $S_{\tilde{\Theta}_{0}}$ and $S_{\tilde{\Theta}_{1}}$ with $n_{1}=n_{0}=2$
$\pi_{A}(Y=y)=\tilde{\pi}_{A}(Y=y\mid Y\in S_{\tilde{\Theta}_{0}})=\tilde{\pi}_{B}(Y=y\mid Y\in S_{\tilde{\Theta}_{1}})=\pi_{B}(Y=y).$
This completes the proof.
Lemma A.2.
Suppose $n$ is even. Let $\tilde{\Theta}_{0}$, $\tilde{\Theta}_{1}$ be as defined in (4)-(5). Define
$C\coloneqq\bigl\{(a,b)\in\mathbb{Z}^{2}:\ 0\leq a\leq n_{1},\ 0\leq b\leq n_{0},\ a+b\equiv n_{1}+1\pmod{2}\bigr\}$ (6)
to be the set of points in $\{0,\ldots,n_{1}\}\times\{0,\ldots,n_{0}\}$ whose sum has the opposite parity of $n_{1}$. Then
$S_{\tilde{\Theta}_{0}}=C=S_{\tilde{\Theta}_{1}}.$
Proof.
First, observe that since $n=n_{0}+n_{1}$ is even, we have that $n_{0}\equiv n_{1}\pmod{2}$.
Step 1: $S_{\tilde{\Theta}_{0}}\subseteq C$. If $(a,b)\in S_{\tilde{\Theta}_{0}}$ then $(a,b)\in S^{Y}_{\theta_{0}(n_{at})}$ for some $n_{at}$ such that $n_{at}\equiv n_{1}+1\pmod{2}$. It is immediate from the expression for $S^{Y}_{\theta_{0}(n_{at})}$ in (2) that $a+b=n_{at}\equiv n_{1}+1\pmod{2}$. Moreover, by definition, $S^{Y}_{\theta_{0}(n_{at})}$ is the support of $Y$ under $\theta_{0}(n_{at})$ and thus $0\leq a\leq n_{1}$ and $0\leq b\leq n_{0}$. Thus, $(a,b)\in C$.
Step 2: $S_{\tilde{\Theta}_{1}}\subseteq C$. If $(a,b)\in S_{\tilde{\Theta}_{1}}$, then $(a,b)\in S^{Y}_{\theta_{1}(n_{c})}$ for some odd $n_{c}$ such that $1\leq n_{c}\leq n-1$. It is immediate from the expression for $S^{Y}_{\theta_{1}(n_{c})}$ in (3) that $b-a=n_{0}-n_{c}$ and hence
$a+b\equiv b-a\equiv n_{0}-n_{c}\pmod{2}$,
where the first $\equiv$ uses the fact that $2a\equiv 0\pmod{2}$. Using the facts that $n_{0}\equiv n_{1}\pmod{2}$ and $n_{c}$ is odd, we see that $n_{0}-n_{c}\equiv n_{0}-1\equiv n_{1}+1\pmod{2}$. Further, by construction $S^{Y}_{\theta_{1}(n_{c})}$ is the support of $Y$ under $\theta_{1}(n_{c})$ and thus $0\leq a\leq n_{1}$ and $0\leq b\leq n_{0}$. Thus, $(a,b)\in C$.
Step 3: $C\subseteq S_{\tilde{\Theta}_{0}}$. Take $(a,b)\in C$ and set $n_{at}\coloneqq a+b$ and $m\coloneqq a$. Then $n_{at}\equiv n_{1}+1\pmod{2}$, and $0\leq n_{at}\leq n_{1}+n_{0}=n$. We claim $(a,b)\in S^{Y}_{\theta_{0}(n_{at})}$. Indeed, $b\leq n_{0}$ implies $n_{at}-n_{0}=a+b-n_{0}\leq a$, so $m=a\geq\max(0,n_{at}-n_{0})$. Also $a\leq n_{1}$ and $a\leq a+b=n_{at}$, so $m=a\leq\min(n_{1},n_{at})$. Thus $(a,b)=(m,n_{at}-m)\in S^{Y}_{\theta_{0}(n_{at})}\subseteq S_{\tilde{\Theta}_{0}}$.
Step 4: $C\subseteq S_{\tilde{\Theta}_{1}}$. Take $(a,b)\in C$ and set
$m\coloneqq a,\qquad n_{c}\coloneqq a+n_{0}-b.$
First, $0\leq n_{c}\leq n$ because $0\leq a\leq n_{1}$ and $0\leq b\leq n_{0}$ and $n_{0}+n_{1}=n$. Second, $n_{c}$ is odd:
$n_{c}=a+n_{0}-b=(a-b)+n_{0}\equiv(a+b)+n_{0}\pmod{2}$,
where the last $\equiv$ uses the fact that $2b\equiv 0\pmod{2}$. Further, since $(a,b)\in C$, $a+b\equiv n_{1}+1\pmod{2}$, so
$n_{c}\equiv(n_{1}+1)+n_{0}\equiv n+1\equiv 1\pmod{2}$,
where we use the fact that $n$ is even. Finally, we check that $m=a$ lies in the admissible range for $S^{Y}_{\theta_{1}(n_{c})}$. Because $n_{c}-n_{0}=a-b$, we have $\max(0,n_{c}-n_{0})=\max(0,a-b)\leq a=m$. Also $m=a\leq n_{1}$ and $m=a\leq n_{c}$ (since $n_{c}=a+n_{0}-b\geq a$). Therefore $(a,b)=(m,m+n_{0}-n_{c})\in S^{Y}_{\theta_{1}(n_{c})}\subseteq S_{\tilde{\Theta}_{1}}$.
Combining Steps 1–4 proves $S_{\tilde{\Theta}_{0}}=C=S_{\tilde{\Theta}_{1}}$. ∎
See 4.2
Proof.
Fix some $n$ that is even. By Lemma˜A.1, there exist priors $\pi_{A}$ and $\pi_{B}$ such that $\pi_{A}(\theta\in\Theta_{0})=1$, $\pi_{B}(\theta\in\Theta_{1})=1$, and $\pi_{A}(Y=y)=\pi_{B}(Y=y)$ for all $y$. Observe that
and hence
It follows that there exists some $\theta_{1}\in\operatorname{supp}(\pi_{B})\subseteq\Theta_{1}$ such that $E_{\theta}[\delta(Y)]\leq\alpha$, which gives the first desired result.
For the second result, fix some odd $n$ and test $\delta$ such that $\sup_{\theta\in\Theta_{0}}E_{\theta}[\delta(Y)]\leq\alpha$. As in the proof to Lemma˜A.1, let $\tilde{\pi}_{A}$ and $\tilde{\pi}_{B}$ be the multinomial priors over $\theta$ with parameters $p=(0.5,0.5,0,0)$ and $p=(0,0,0.5,0.5)$, respectively. As argued in the proof to Lemma˜A.1, $\tilde{\pi}_{A}(Y=y)=\tilde{\pi}_{B}(Y=y)$ for all $y$. This implies that
$E_{\theta\sim\tilde{\pi}_{B}}[E_{\theta}[\delta(Y)]]=E_{\theta\sim\tilde{\pi}_{A}}[E_{\theta}[\delta(Y)]]\leq\alpha,$
where the inequality uses the fact that, by construction, $\tilde{\pi}_{A}(\theta\in\Theta_{0})=1$ and that $\delta$ controls size. Next, observe that $\tilde{\pi}_{B}(\theta\in\Theta_{0})=\tilde{\pi}_{B}(\theta\in\{(0,0,n,0),(0,0,0,n)\}=0.5^{n-1}=:\epsilon$. By iterated expectations, we then have
$E_{\theta\sim\tilde{\pi}_{B}}[E_{\theta}[\delta(Y)]]=\epsilon E_{\theta\sim\tilde{\pi}_{B}}[E_{\theta}[\delta(Y)]\mid\theta\in\Theta_{0}]+(1-\epsilon)E_{\theta\sim\tilde{\pi}_{B}}[E_{\theta}[\delta(Y)]\mid\theta\in\Theta_{1}]$
and hence
$E_{\theta\sim\tilde{\pi}_{B}}[E_{\theta}[\delta(Y)]\mid\theta\in\Theta_{1}]$ $=\frac{1}{1-\epsilon}E_{\theta\sim\tilde{\pi}_{B}}[E_{\theta}[\delta(Y)]]-\frac{\epsilon}{1-\epsilon}E_{\theta\sim\tilde{\pi}_{B}}[E_{\theta}[\delta(Y)]\mid\theta\in\Theta_{0}]$
$\leq\frac{\alpha}{1-\epsilon}=\alpha\frac{1}{1-2^{-(n-1)}}=\alpha\left(1+\frac{1}{2^{n-1}-1}\right)$
where the inequality uses (A) together with the fact that $\delta(Y)\geq 0$. The result then follows from the fact that $\inf_{\theta\in\Theta_{1}}E_{\theta}[\delta(Y)]\leq E_{\theta\sim\tilde{\pi}_{B}}[E_{\theta}[\delta(Y)]\mid\theta\in\Theta_{1}]$. ∎
@@
Let $\pi(t;\theta_{1})$ be the PMF for $\mathrm{Multinomial}(n,\theta_{1}/n)$. Then
$w_{n}(t;\theta_{1})=\frac{\pi(t;\theta_{1})}{\sum_{t\in\Theta_{1}}\pi(t;\theta_{1})}\mathbbm{1}(t\in\Theta_{1}).$
Let $q=\theta_{1}/n$ and observe that $g_{q}(y)=\sum_{t\in\Theta_{0}\cup\Theta_{1}}f_{t}(y)\pi(t;\theta_{1})$. By Proposition˜3.1, there exists some $p\in\mathcal{P}_{0}$ such that $g_{q}(y)=g_{p}(y)$. We thus have the following expansion of WAP,
$\mathrm{WAP}(\delta;\theta_{1})$ $=\sum_{t\in\Theta_{0}\cup\Theta_{1}}\sum_{y}\delta(y)f_{t}(y)\frac{\pi(t;\theta_{1})}{\sum_{t\in\Theta_{1}}\pi(t;\theta_{1})}\mathbbm{1}(t\in\Theta_{1})$
$=\underbrace{\left({\sum_{t\in\Theta_{1}}\pi(t;\theta_{1})}\right)^{-1}}_{=:Z_{1}^{-1}}\sum_{y}\delta(y)\left[{\sum_{t\in\Theta_{0}\cup\Theta_{1}}f_{t}(y)\pi(t;\theta_{1})-\sum_{t\in\Theta_{0}}f_{t}(y)\pi(t;\theta_{1})}\right]$
$=\frac{1}{Z_{1}}\left[{E_{g_{q}}[\delta(Y)]-\sum_{t\in\Theta_{0}}E_{t}[\delta(Y)]\pi(t;\theta_{1})}\right]$
@@
where recall $v=\frac{1}{n}\min\{\theta_{1,c},\theta_{1,d}\}=\min\{q_{c},q_{d}\}$. This proves the first inequality. The second inequality follows from examining that $v\geq 1/n$ and that
@@
Let $\delta_{0}(Y)=\alpha\cdot 1\{(Y_{T},Y_{U})\not\in\{(n_{1},0),(0,n_{0})\}\}$ be the test that rejects with probability 0 if $Y$ is either $(n_{1},0)$ or $(0,n_{0})$, and rejects with probability $\alpha$ otherwise. Let $S^{Y}_{\theta}$ be the support of $Y$ under parameter $\theta$. Note that by construction, for any $\theta_{0}=(n_{at},n_{nt},n_{d},n_{c})\in\Theta_{0}$, we have that
@@
$E_{\theta_{0}}[\delta_{0}(Y)]\leq\alpha\text{ if }E_{\theta_{0}}[\delta_{1}(Y)]=0.$
@@
it suffices to show that, for all $\theta_{1}\in\Theta_{1}$,
@@
Finally, we show that there exists $\theta_{1}$ such that $E_{\theta_{1}}[\delta(Y)]>\alpha$. Since we showed above that $E_{\theta_{1}}[\delta_{0}(Y)]=\alpha$ for all $\theta_{1}\in\Theta_{1}$, it suffices to show that there exists some $\theta_{1}\in\Theta_{1}$ such that $E_{\theta_{1}}[\delta_{1}(Y)]>0$, or equivalently $P_{\theta_{1}}(Y=(n_{1}-1,1))>0$. However, if $\theta_{1}$ corresponds to $(n_{at},n_{nt},n_{d},n_{c})=(0,0,2,n-2)$, then $Y=(n_{1}-1,1)$ obtains when one defier is assigned to treatment and the other to control. ∎
@@
$E_{\theta}[\delta(Y)]\leq\alpha(1+3.125n^{1.5}(1-\epsilon)^{n})\to\alpha$
as $n\to\infty$. ∎
@@
$\pi(\Theta_{0}\mid Y=y)=\pi(\theta_{0}\mid Y=y)=1-\pi(\theta_{1}\mid Y=y)=\frac{f_{\theta_{0}}(y)}{f_{\theta_{1}}(y)+f_{\theta_{0}}(y)}.$
Since $f_{\theta_{0}}\neq f_{\theta_{1}}$ for some $y$ in the support of one of $f_{\theta_{0}}(\cdot)$ and $f_{\theta_{1}}(\cdot)$, $\pi(\theta_{0}\mid Y=y)\neq 0.5=\pi(\theta_{0})$ with positive probability. This concludes the proof of the first part.
@@
See 5.1
Proof.
Observe that for any prior $\pi$ over $\theta$,
$\inf_{c(\cdot)}\sup_{\theta}R(c,\theta)$ $\geq\inf_{c(\cdot)}E_{\theta\sim\pi}[R(c,\theta)].$
For a (possibly randomized) classifier $c(\cdot)$, it follows that:
$E_{\theta\sim\pi}[R(c,\theta)]$ $=E_{\theta\sim\pi}[E_{Y\sim\pi_{Y\mid\theta}}[c(Y)\ell(1,\theta)+(1-c(Y)\ell(0,\theta)]]$
$=E_{Y\sim\pi_{Y}}[E_{\theta\sim\pi_{\theta\mid Y}}[c(Y)\ell(1,\theta)+(1-c(Y)\ell(0,\theta)]]$
$=E_{Y\sim\pi_{Y}}\Big[(1-\pi(\theta\in\Theta_{0}\mid Y))\,c(Y)+\pi(\theta\in\Theta_{0}\mid Y)\,(1-c(Y))\Big],$
where for clarity in the previous display, we use e.g. $\pi_{Y}$ to denote the distribution of $Y$ induced by the prior $\pi$, and we make explicit the measures over which random variables are taken.
By Proposition˜5.1 part 2, there exists a prior $\pi$ such that $\pi(\theta\in\Theta_{0}\mid Y)=\pi(\theta\in\Theta_{0})=0.5$ ($\pi$-a.s.), in which case we have
$E_{\theta\sim\pi}[R(c,\theta)]$ $=E_{Y\sim\pi_{Y}}\Big[\tfrac{1}{2}c(Y)+\tfrac{1}{2}(1-c(Y))\Big]=\tfrac{1}{2}.$
Hence for every classifier $c(\cdot)$, its Bayes risk under $\pi$ equals $1/2$, so in particular
$\inf_{c(\cdot)}E_{\theta\sim\pi}[R(c,\theta)]=\tfrac{1}{2},$
and therefore
$\inf_{c(\cdot)}\sup_{\theta}R(c,\theta)\geq\tfrac{1}{2}.$
On the other hand, the trivial randomized classifier, $c(y)=1/2$ for all $y$, has
$R(\theta,c)=\tfrac{1}{2}\quad\text{for all }\theta,$
and thus
$\inf_{c(\cdot)}\sup_{\theta}R(c,\theta)\leq\tfrac{1}{2}.$
Combining the two inequalities yields $\inf_{c(\cdot)}\sup_{\theta}R(c,\theta)=1/2$. ∎
Proposition A.1.
Suppose $n$ is odd. For any $c\in(0,1)$, there exists a prior $\pi$ such that $\pi(\theta\in\Theta_{0})=c$ and
$E_{Y\sim\pi_{Y}}\left[\left|\pi(\theta\in\Theta_{0}\mid Y)-\pi(\theta\in\Theta_{0})\right|\right]=2^{-(n-2)}(1-c),$
where $\pi_{Y}$ is the push-forward of $Y$ under $\theta\sim\pi$.
Proof.
As in the proof to Lemma˜A.1, let $\tilde{\pi}_{A}$ and $\tilde{\pi}_{B}$ be the multinomial priors over $\theta$ with parameters $p=(0.5,0.5,0,0)$ and $p=(0,0,0.5,0.5)$, respectively. As argued in the proof to Lemma˜A.1, $\tilde{\pi}_{A}(Y=y)=\tilde{\pi}_{B}(Y=y)$ for all $y$. Observe that $\tilde{\pi}_{A}(\theta\in\Theta_{0})=1$ and hence $\tilde{\pi}_{A}(\theta\in\Theta_{0}\mid Y)=1$. Next, observe that $\tilde{\pi}_{B}(\theta\in\Theta_{0})=\tilde{\pi}_{B}(\theta\in\{(0,0,n,0),(0,0,0,n)\}=0.5^{n-1}=:\epsilon$. Let $C:=\{(n_{1},0),(0,n_{0})\}$ denote the corners of the support of $Y$, and observe that under $\tilde{\pi}_{B}$, $Y\in C$ if and only if $\theta\in\Theta_{0}$: that is, $\tilde{\pi}_{B}(Y\in C\mid\theta\in\Theta_{0})=1$, $\tilde{\pi}_{B}(Y\in C\mid\theta\in\Theta_{1})=0$. It follows that $\tilde{\pi}_{B}(Y\in C)=\epsilon$, and by Bayes’ rule, $\tilde{\pi}_{B}(\theta\in\Theta_{0}\mid Y)=1\{Y\in C\}$.
Now, let $\pi(\theta)=\omega\tilde{\pi}_{A}(\theta)+(1-\omega)\tilde{\pi}_{B}(\theta)$ be the mixture of $\tilde{\pi}_{A}$ and $\tilde{\pi}_{B}$, with the mixture weight $\omega:=\frac{c-\epsilon}{1-\epsilon}$ chosen so that $\pi(\theta\in\Theta_{0})=c$. Observe that
$\pi(\theta\in\Theta_{0}\mid Y)$ $=\frac{\pi(\theta\in\Theta_{0},Y)}{\pi(Y)}$
$=\frac{\omega\tilde{\pi}_{A}(\theta\in\Theta_{0}\mid Y)\tilde{\pi}_{A}(Y)+(1-\omega)\tilde{\pi}_{B}(\theta\in\Theta_{0}\mid Y)\tilde{\pi}_{B}(Y)}{\pi(Y)}$
$=\omega\tilde{\pi}_{A}(\theta\in\Theta_{0}\mid Y)+(1-\omega)\tilde{\pi}_{B}(\theta\in\Theta_{0}\mid Y)$
where the first line uses the definition of conditional probability, the second line uses the definition of $\pi$, and the third line uses the fact that $\tilde{\pi}_{A}(Y)=\tilde{\pi}_{B}(Y)=\pi(Y)$. It follows from the previous display and our earlier derivations of the posterior probabilities that
$\pi(\theta\in\Theta_{0}\mid Y)=\omega+(1-\omega)1\{Y\in C\}.$
We thus see that
$E_{Y\sim\pi_{Y}}\left[\left|\pi(\theta\in\Theta_{0}\mid Y)-\pi(\theta\in\Theta_{0})\right|\right]$ $=\pi(Y\in C)(1-c)+(1-\pi(Y\in C))(c-\omega)$
$=\epsilon\left(1-c\right)+(1-\epsilon)(c-\omega)$
$=2\epsilon(1-c).$
The result then follows from the definition of $\epsilon=2^{-(n-1)}$. ∎
Corollary A.1.
Suppose $n$ is odd. Then $\inf_{c(\cdot)}\sup_{\theta}R(c,\theta)\geq 0.5-2^{-(n-1)}$.
Proof.
As argued in the proof to Corollary˜5.1, for any prior $\pi$, $\inf_{c(\cdot)}\sup_{\theta}R(c,\theta)\geq\inf_{c(\cdot)}E_{\theta\sim\pi}[R(c,\theta)]$ and
$E_{\theta\sim\pi}[R(c,\theta)]=E_{Y\sim\pi_{Y}}\Big[(1-\pi(\theta\in\Theta_{0}\mid Y))\,c(Y)+\pi(\theta\in\Theta_{0}\mid Y)\,(1-c(Y)).$
From Proposition˜A.1, there exists $\pi$ such that $\pi(\theta\in\Theta_{0})=0.5$ and $E_{Y\sim\pi_{Y}}[|\pi(\theta\in\Theta_{0}\mid Y)-\pi(\theta\in\Theta_{0})|]=2^{-(n-1)}$. Adding and subtracting $E_{Y\sim\pi_{Y}}[(1-\pi(\Theta_{0}))c(Y)+\pi(\Theta_{0})(1-c(Y))]=0.5$ from (A), we see that
$E_{\theta\sim\pi}[R(c,\theta)]$ $=0.5+E_{Y\sim\pi_{Y}}\Big[(\pi(\Theta_{0}\mid Y)-\pi(\Theta_{0}))(1-2c(Y))\Big]$
$\geq 0.5-E_{Y\sim\pi_{Y}}[|\pi(\Theta_{0}\mid Y)-\pi(\Theta_{0})|\cdot|1-2c(Y)|]$
$\geq 0.5-E_{Y\sim\pi_{Y}}[|\pi(\Theta_{0}\mid Y)-\pi(\Theta_{0})|]$
$=0.5-2^{-(n-1)}$
where the second line uses the triangle inequality, and the third uses the fact that $c(Y)\in[0,1]$. This completes the proof. ∎